• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 23:39:10 on September 26,1999

    Entry number 27236

    keyword=shift summary - 9/26/1999 swing shift

    Finish up 53 degree cm point. Lose ~1 hour dur to accelerator problems
    + shift checks.

    Because e-arm change is 5 minutes to setup for 12C data, but it also
    is 5 minutes to change targets, decide to do 12C data to start the 70 deg cm point.

    Run 1854: start 17:40, 20 minutes long, 16 k e+12C events for alignment.

    Switch to radiator out, raster on, proton target.

    Run 1855: start 18:10, end 19:10, 101 k events

    Beam current oscillations have reappeared, as of ~ 18:00. Notice this around
    18:30, warn MCC around 18:50.

    18:25: Notice filtering ended at run 1836, taken about 3 pm - 5 pm yesterday.
    Restart it up. It looks like no disk has > 15% free space, so it is just marking time
    for now.

    18:35 shift checklist done.

    18:53 MCC says accelerator energy stepped off by 10^-3 for 1 minute
    starting at this time, then they brought it back.

    19:10 switch to e + deuteron target
    Run 1856 started 18:15

    19:22 Coda 100% dead. All components seem to be up. Called Bob.

    19:30 Beam is down for 5 minutes for injector correction.

    19:36 Beam is back.

    19:47 Bob finished remote diagnostic (see log entry 27231). Ended run1856.
    Started run 1857. Stopped 21:20 with 259 k events.

    20:04 MCC called. They have someone looking into observed since 17:30 current
    oscillations (amplitude of oscillations ~5uA, period ~20min). They suspect something with RF.
    20:45 Oscillations come from injector MCC says. Experts coming in.

    21:21 Started run 1858. Beam taken away 21:45 for injector work. Run stopped 22:10 with 95 k events.

    22:45 MCC calls, phase problem in injector still being tracked down, then
    injector person needed for retune. Experts estimate 1 hour + ==> beam
    maybe midnight. Decide to put in radiator for benefit of overnight shift
    not having to do much.

    Since this shift saw the start os current oscillations and many hours
    without beam, + 1 DAQ crash, it has not been as positive an experience
    as one would have hoped. I submit this early, assuming beam will not return
    in next 15 minutes.